OpenStudy (anonymous):

The midpoint of line UV is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (binary3i):

7,-27

OpenStudy (anonymous):

( x1 + x2)/2 = x - midpoint So (3 + x)/2 = 5 Times by 2 3 + x = 10 or x= 7 Can you do the same for y?

OpenStudy (hoblos):

(7,-27)

OpenStudy (amistre64):

if we have the end point at (0,0) then the midpoint is easy to work with :) (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) since the midpoint is half way, double it to get to the end. (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) 4 -32 --------- (x-1 , y-37) we modified (x=3,y=5) to begin with so lets sub those values back in. (3-1, 5-37) = (2, -27)

OpenStudy (amistre64):

or simply, midpoint - endpoint; double it, and add back in the endpoint